# GABAR: Graph Attention-Based Action Ranking for Relational Policy Learning

### v2

@@ Abstract
We propose a novel approach to learn relational policies for classical planning based on learning to rank actions. We introduce a new graph representation that explicitly captures action information and propose a Graph Neural Network architecture augmented with Gated Recurrent Units (GRUs) to learn action rankings. Our model is trained on small problem instances and generalizes to significantly larger instances where traditional planning becomes computationally expensive. Experimental results across standard planning benchmarks demonstrate that our action-ranking approach achieves generalization to significantly larger problems than those used in training.
Introduction
Classical planning tackles the problem of finding action sequences to achieve goals in deterministic environments. While traditional planners can solve small problems optimally using search and heuristics, they often struggle with scalability. This has motivated research into learning general relational policies from small solved problems, which can then be applied to significantly larger instances (Ståhlberg, Bonet, and Geffner 2022a). The key insight behind learning-based planning is that optimal solutions to small problems often reveal patterns that generalize to bigger problems. For example, in Blocks World, solutions to $4$ block problems can teach a policy to stack blocks bottom-up that generalizes to problems with $20$ or more blocks. The advantage of relational learning is its ability to capture compositional structure. This compositionality also enables strong generalization (Zambaldi et al. 2018), (Fern, Yoon, and Givan 2003) (Džeroski, De Raedt, and Driessens 2001).
@@ Introduction
Recently, several neural approaches have been proposed for learning domain-specific heuristics. (Toyer et al. 2020) learn neural heuristics that generalize across problem instances but still rely on search for planning. Similarly, (Shen, Trevizan, and Thiébaux 2020) use hypergraph networks to learn heuristics that can transfer across problems. Other approaches that learn value functions for planning include (Chen, Thiébaux, and Trevizan 2023), (Chen, Thiébaux, and Trevizan 2024). While these approaches show promise in learning useful search guidance, they suffer from the inherent computational overhead of the search process during planning.
An alternative line of work focuses on learning policies that can directly select actions without search. Value-based methods like (Ståhlberg, Bonet, and Geffner 2022b), (Ståhlberg, Bonet, and Geffner 2022a) learn a value function that induces a greedy policy by selecting actions leading to states with minimum estimated cost-to-go. While effective, these approaches face two key limitations: First, value functions are not only quite complex and challenging to learn, but also are unnecessary for action selection. Second, in domains where optimal planning is NP-hard (like the Blocks World), optimal value functions are not easily generalizable to larger problem sizes (Gupta and Nau 1992).
@@
Learning to Rank approaches have shown promise in planning domains but have primarily focused on ranking states rather than actions. (Garrett, Kaelbling, and Lozano-Pérez 2016) pioneered this direction by using RankSVM to learn state rankings using hand-crafted features. More recently, (Chrestien et al. 2024) and (Hao et al. 2024) demonstrated that learning to rank states can be more effective than learning precise heuristic values, as the relative ordering of states is sufficient for guiding search. However, these approaches still fundamentally rely on search during execution, inheriting signficant computational overhead.
@@
Our work is inspired by the effectiveness of graph neural networks (GNNs) to represent and learn general relational policies such as (Ståhlberg, Bonet, and Geffner 2022b, 2023, 2024; Chen, Thiébaux, and Trevizan 2024). We introduce a novel architecture GABAR (Graph Attention-Based Action Ranking), which directly learns to rank actions rather than estimating value functions. GABAR introduces (1) an action-centric graph representation that explicitly captures how objects participate in actions and (2) a GNN architecture with Gated Recurrent Units (GRUs) that learns how object representations update based on their role in actions. One of our key insights is that ranking actions that are applicable in the same state often turns out to be easier and more generalizable than ranking states by their distances to the goals. Through experiments on standard benchmarks, we show GABAR achieves generalization to significantly larger problems than those used for training.
Problem Setup
@@
Graph Attention-Based Action Ranking
Figure 1: GABAR’s architecture for action extraction. (a) Graph representation: The input PDDL problem is converted into a graph with four types of nodes (predicate, object, action schema, and global) connected by predicate-object and action-object edges that encode state and grounded action information. (b) GNN encoder: Processes the graph through $L$ rounds of message passing where edge, node, and global representations are sequentially updated based on their interactions. (c) Action decoder: Uses the final global embedding to construct a grounded action through a GRU-based decoder sequentially - first selecting an action schema, then iteratively choosing objects for each parameter position until a complete grounded action is formed.
@@
Learning general policies for classical planning domains requires effectively processing variable-sized states and selecting appropriate actions that generalize across problem instances. We introduce GABAR (Graph Attention-based Decoder for Action Ranking), a novel architecture that directly learns to rank and select actions rather than learning value functions. GABAR consists of four key components: (1) a graph-based state representation that explicitly captures grounded action information, (2) a neural encoder that processes this rich representation, and (3) a GRU-based decoder that sequentially constructs complete actions. This section details each component and explains how they work together to enable effective action selection.
System Overview
@@
Given a planning instance, I in PDDL format (Fox and Long 2003), along with the set of ground actions, GABAR operates by first converting this to a graph structure that makes explicit the relationships between objects, predicates, and potential actions. This graph is then processed through our neural architecture to rank actions(as described in Fig 1). Then, the highest-ranked applicable action is executed to reach the next state I’. This process is repeated until a goal state is reached. To ensure the execution terminates, the system maintains a history of visited states and avoids actions that would lead to previously visited states. The execution continues until either reaching a goal state or terminating if no unvisited successor states are available or if the maximum execution length (1000 in our experiments) is exceeded.
Graph Representation
@@
We introduce a novel graph representation for classical planning tasks that captures the structural relationships between objects, predicates, and actions and the semantic information needed for learning action ranking effectively. Our representation $G=(V,E,X,R)$ consists of a set of nodes $V$, edges $E$, node features $X$, and edge features $R$.
@@
The node set $V=O\cup P\cup A\cup\{g\}$ where $O$, $P$,and $A$ represent sets of domain objects, grounded predicates, and action schemas respectively and $g$ is a global node that aggregates graph-level information. The edge set $E=E_{\text{pred}}\cup E_{\text{act}}$, where $E_{\text{pred}}$ is the set of edges between predicates and their argument objects and similarly $E_{\text{act}}$ is the set of edges between action schemas and their argument objects.
Node Features
@@
The node feature function $X:V\rightarrow\mathbb{R}^{d}$ maps each node to a feature vector that encodes type and semantic information. The feature vector is constructed by concatenating several one-hot encoded segments:
For any node $v\in V$: $X(v)=[X_{\text{type}}(v)\parallel X_{\text{act}}(v)\parallel X_{\text{pred}}(v)\parallel X_{\text{obj}}(v)]$, where:
@@
  • $X_{\text{pred}}\in\{0,1\}^{2|P|}$: Encoding for predicates, where first $|P|$ bits indicate predicate type and next $|P|$ bits indicate goal predicates ((if $v$ is a predicate node))
  • $X_{\text{obj}}\in\{0,1\}^{|T|}$: One-hot encoding of object type (if $v$ is an object node), where $T$ is the set of object types
Edge Features
The edge feature function $R:E\rightarrow\mathbb{R}^{k}$ maps each edge to a feature vector encoding edge type and role information:
For any edge $e\in E$: $R(e)=[R_{\text{type}}(e)\parallel R_{\text{pred}}(e)\parallel R_{\text{act}}(e)]$, where:
@@
    • Binary vector indicating which predicates are satisfied by the object in the grounded action ($|P|$ bits)
Global Features
The global node $g$ is initialized with a zero vector in $\mathbb{R}^{h}$ where $h$ is the chosen hidden representation dimension. This node can be used to aggregate and propagate graph-level information during message passing.
@@
Our neural architecture processes this graph representation through multiple components designed to handle the challenges of processing variable-sized inputs, capturing long-range dependencies between objects and actions, and making sequential decisions to construct complete actions. We detail each component below:
Graph Neural Network Encoder
@@
Graph Neural Networks (GNNs) (Scarselli et al. 2008) are particularly well-suited for encoding states in planning problems as they can naturally process relational structures while being invariant to permutations and handling varying input sizes. This allows them to learn patterns that generalize across different problem instances within the same domain, regardless of the number of objects involved. The key insight is that planning states are inherently relational - objects interact through predicates and actions - and GNNs can capture these relationships through message passing between nodes and edges.
Our GNN encoder processes the input graph through $L$ layers of message passing between nodes, edges, and a global node. At each layer $l$, the updates proceed as follows:
First, edge embeddings are updated based on their incident nodes and the global context:
$\mathbf{e}^{l+1}_{ij}=\phi_{e}([\mathbf{e}^{l}_{ij};\mathbf{v}^{l}_{i};\mathbf{v}^{l}_{j};\mathbf{g}^{l}])$ (1)
where $\mathbf{e}^{l}_{ij}$ is the embedding of edge $(i,j)$ at layer $l$, $\mathbf{v}^{l}_{i}$ and $\mathbf{v}^{l}_{j}$ are the embeddings of its incident nodes, $\mathbf{g}^{l}$ is the global node embedding, [;][;][ ; ] denotes concatenation, and $\phi_{e}$ is a learnable neural network.
Next, node embeddings are updated using information from their connected edges and the global context:
$\mathbf{v}^{l+1}_{i}=\phi_{v}([\mathbf{v}^{l}_{i};\text{AGG}(\{\mathbf{e}^{l+1}_{ij}|j\in\mathcal{N}(i)\});\mathbf{g}^{l}])$ (2)
Where AGG is a permutation-invariant aggregation function (we use attention weighted sum), $\mathcal{N}(i)$ denotes the neighbors of node $i$, and $\phi_{h}$ is a learnable network.
Finally, the global node is updated by aggregating information from all nodes and edges:
$\mathbf{g}^{l+1}=\phi_{g}([\mathbf{g}^{l};\text{AGG}(\{\mathbf{v}^{l+1}_{i}|i\in\mathcal{V}\});\text{AGG}(\{\mathbf{e}^{l+1}_{ij}|(i,j)\in\mathcal{E}\})])$ (3)
where $\phi_{g}$ is a learnable network that combines the previous global state with aggregated node and edge information.
@@
This architecture makes several key design choices motivated by the planning domain. First, we explicitly model and update edge representations because edges in our graph capture crucial information about action applicability. This edge information helps guide the model toward selecting valid and effective actions. Second, we include a global node that can rapidly propagate information across the graph. This is particularly important as planning problems scale up - without the global node, information would need to flow through many message-passing steps to reach distant parts of the graph. The global node acts as a shortcut, allowing the model to maintain a comprehensive view of the planning state even as the number of objects and relations grows. After $L$ rounds of message passing have been performed, the final global node embedding $g^{l+1}$ captures the relevant planning context needed for action selection.
GRU-based Action Decoder
One of the key novelties of our architecture is a GRU-based action decoder that supports actions with an arbitrary number of parameters. The decoder transforms the final global graph representation $g_{l+1}$ into a sequence of parameter decisions that fully specify an action. This sequential process is managed by a Gated Recurrent Unit (GRU) that maintains a hidden state $h_{i}$ representing the partial action construction:
$v_{1}$ $=\text{GRU}(g_{v+1},0)$ // Initialize with global state
@@
The GRU architecture allows the decoder to maintain relevant context while making a sequence of interdependent decisions. This is crucial for ensuring that selected objects form valid and effective action instantiations.
During training, action selection proceeds greedily in two phases: First, the decoder selects an action schema using the initial hidden state:
$\text{score}(a)=\text{MLP}([v_{1};a])\quad\text{for each action schema }a$
$a^{*}=\arg\max_{a\in A}\text{softmax}(\text{score}(a))$
Then, it iteratively selects the action parameters:
$\text{score}(o)=\text{MLP}([v_{i};o])\quad\text{for each candidate object }o$
$o^{*}=\arg\max_{o\in O}\text{softmax}(\text{score}(o))$
$v_{i+1}=\text{GRU}(v_{i},\text{embed}(o^{*}))$
Since greedy parameter selection is often too myopic, we employ beam search to explore multiple choices in parallel. For a beam width $k$, at each step, we maintain the $k$ highest-scoring partial sequences. The final output is a ranked list of $k$ action groundings $(a,o_{1},\ldots,o_{n})$ along with their accumulated scores. This aligns with our goal of learning to rank actions rather than just selecting individual best choices - while training optimizes for selecting the optimal action, the learned model provides a ranking of top $k$ grounded actions during execution. The planner can then use this ranking to make more informed decisions, such as incorporating additional criteria such as cycle avoidance.
Data Generation and Training

### v3

# Graph Neural Network Based Action Ranking for Planning

@@ Abstract
We propose a novel approach to learn relational policies for classical planning based on learning to rank actions. We introduce a new graph representation that explicitly captures action information and propose a Graph Neural Network (GNN) architecture augmented with Gated Recurrent Units (GRUs) to learn action rankings. Unlike value-function based approaches that must learn a globally consistent function, our action ranking method only needs to learn locally consistent ranking, which is more sample-efficient. Our model is trained on data generated from small problem instances that are easily solved by planners and is applied to significantly larger instances where planning is computationally prohibitive. Experimental results across standard planning benchmarks demonstrate that our action-ranking approach not only achieves better generalization to larger problems than those used in training but also outperforms multiple baseline (value function and action ranking) methods in terms of success rate and plan quality.
1 Introduction
Classical planning tackles the problem of finding action sequences to achieve goals in deterministic environments. While traditional planners can solve small problems optimally using search and heuristics, they often struggle with scalability. This has motivated research into learning general relational policies from small solved problems, which can then be applied to significantly larger instances (expressive_v1, ). The key insight behind learning-based planning is that optimal solutions to small problems often reveal patterns that generalize to bigger problems. For example, in Blocks World, solutions to $4$ block problems can teach a policy to stack blocks bottom-up that generalizes to problems with $20$ or more blocks. The advantage of relational learning is its ability to capture compositional structure, which in turn enable strong generalization (relational_drl, ), (approx_Policy_iter, ) (relational_rl, ) relational_mdp_learning .
@@ 1 Introduction
Recently, several neural approaches have been proposed in the planning domain. Some methods learn value functions to guide search processes (strips_hgn, ; goose_v2, ), while others learn value functions that induce greedy policies by selecting actions leading to states with minimum estimated cost-to-go (unsupervised_gnn, ; expressive_v1, ). While these approaches show promise, they all face a fundamental challenge: value functions are not only complex and challenging to learn, but since optimal planning is NP-hard in most domains (complexity-planning, ; blocks_np_hard, ), there is no reason to think that optimal value functions generalize to larger problem sizes. In our work, instead of trying to learn optimal value functions we focus on learning a policy that ranks good actions higher in any given state.
@@
Learning to rank approaches have shown promise in planning domains but have primarily focused on ranking states rather than actions. (ranking_og, ) pioneered this direction by using RankSVM to learn state rankings using hand-crafted features. More recently, (chrestien_optimize, ), and gbfs_v2 demonstrated that learning to rank states can be more effective than learning precise heuristic values, as the relative ordering of states is sufficient for guiding search. However, these approaches still fundamentally rely on search during execution, inheriting significant computational overhead. The other drawback is that, while ranking states is easier than learning the exact value function, they are still trying to learn a globally consistent state ranking function which is challenging. We overcome this issue in our work by learning only local ranking over actions in any given state instead of learning a ranking among all possible states. Unlike other action ranking approaches such as (karia2021learning, ), (rddl_gnn, ), (sr_drl, ), and (policy_gradients, ), we excitability represent action information in our input state representation.
Our work is inspired by the effectiveness of graph neural networks (GNNs) (gnn_base, ) to represent and learn general relational policies such as (unsupervised_gnn, ; policy_gradients, ; policy_beyond_c2, ; goose_v2, ; relational_drl, ) where GNNs have been shown to be able to learn well over state representations for planning problems. GNNs are especially suited for these types of problems as they can handle inputs of varied sizes and can learn from graphs of small size and generalize to larger graphs hamilton2020graph . This is the property we are looking to exploit - train a system on small sized planning problems where it is easy to run planners and collect data and use this learned model to solve larger problems that planners are too slow to solve while simplifying the learning problem by only learning local action rankings.
@@
To achieve this, we introduce a novel architecture GABAR (GrAph neural network Based Action Ranking), which directly learns to rank actions rather than estimating value functions. GABAR consists of three key components: (1) an action-centric graph representation of state that explicitly captures how objects participate in actions (2) a GNN encoder that processes this rich representation, and (3) a GRU-based decoder that sequentially constructs parameterized actions. Our key insight is that ranking actions that are applicable in the same state often turns out to be easier and more generalizable than ranking states by their distances to the goals. Through experiments on standard benchmarks, we show that (a) GABAR achieves generalization to significantly larger problems than those used for training, and (b) it does markedly better on the larger problems when compared to value-function-based methods or other methods that do not include action information in the graph.
2 Problem Setup
@@
3 GNN Based Action Ranking
@@
Learning general policies for classical planning domains requires learning to select appropriate actions effectively to variable-sized states and generalize across problem instances. This section details each component of GABAR, the graph representation, the GNN encoder and the GRU decoder in detail and explains how they work together to enable effective action selection.
3.1 System Overview
@@
Given a planning instance, I in PDDL format pddl , along with the set of ground actions, GABAR operates by first converting this to a graph structure that makes explicit the relationships between objects, predicates, and potential actions. This graph is then processed through our neural architecture to rank actions(as described in Fig 1). Then, the highest-ranked applicable action is executed to reach the next state I’, and the process repeats. To ensure that the execution terminates, the system maintains a history of visited states and avoids actions that would lead to previously visited states. The execution continues until either a goal state is reached or a state is reached with no unvisited successor states, or the maximum execution length (1000 in our experiments) is exceeded.
Figure 1: GABAR’s architecture for action extraction. (a) Graph representation: The input PDDL problem is converted into a graph with four types of nodes (predicate, object, action schema, and global) connected by predicate-object and action-object edges that encode state and grounded action information. (b) GNN encoder: Processes the graph through $L$ rounds of message passing where edge, node, and global representations are sequentially updated (c) Action decoder: Uses the final global embedding to construct a grounded action through a GRU-based decoder sequentially - first selecting an action schema, then iteratively choosing objects for each parameter position until a complete grounded action is formed.
3.2 Graph Representation
@@
We introduce a novel graph representation (shown in Fig 2) for classical planning tasks that captures the structural relationships between objects, predicates, actions, and the semantic information needed to effectively learn action ranking. Our representation $G=(V,E,X,R)$ consists of a set of nodes $V$, edges $E$, node features $X$ and edge features $R$.
@@
The node set $V=O\cup P\cup A\cup\{g\}$ where $O$, $P$,and AA represent sets of domain objects, grounded predicates (predicates instantiated in current and goal states), and action schemas, respectively, and $g$ is a global node that aggregates graph-level information.
The edge set $E=E_{\text{pred}}\cup E_{\text{act}}$, where
$E_{\text{pred}}$ is the set of edges between predicates and their argument objects and similarly $E_{\text{act}}$ is the set of edges between action schemas and their argument objects.
@@
Node Features: The node feature function $X:V\rightarrow\mathbb{R}^{d}$ maps each node to a feature vector that encodes type and semantic information. The feature vector is constructed by concatenating several one-hot encoded segments. For any node $v\in V$, $X(v)=[X_{\text{type}}(v)\parallel X_{\text{act}}(v)\parallel X_{\text{pred}}(v)\parallel X_{\text{obj}}(v)]$, where,
Figure 2: Example graph construction for a simplified blocksworld problem with only onon and clearclear predicates. The left side shows the starting and goal states. The $O$ nodes are the object nodes. In the start state, $O3$ is on $O2$ which is on the table, and $O1$ is on the table. In the goal state, $O3$ is on $O1$ which is on the table, and $O2$ is on the table. The right side shows the constructed graph with action nodes (blue), object nodes (yellow), and predicate nodes (red). The “Pick-up” action connects to object $O1$, while the “Unstack” action connects to objects $O2$ and $O3$. Predicate nodes show the current state (“Clear” for $O1$ and $O3$, “On (O3,O2)”) and goal state (“G-ON(O3,O1)”). Blue edges represent action-object connections, while red edges represent predicate-object relationships.
@@
  • $X_{\text{pred}}\in\{0,1\}^{2|P|}$: Encoding for predicates, where first $|P|$ bits indicate predicate type and next $|P|$ bits indicate goal predicates (if $v$ is a predicate node)
  • $X_{\text{obj}}\in\{0,1\}^{|T|}$: One-hot encoding of the object type (if $v$ is an object node). $T$ is the set of object types
Edge Features: The edge feature function $R:E\rightarrow\mathbb{R}^{k}$ maps each edge to a feature vector that encodes the type of edge and the information about the role. For any edge $e\in E$, $R(e)=[R_{\text{type}}(e)\parallel R_{\text{pred}}(e)\parallel R_{\text{act}}(e)]$, where,
@@
    • Binary vector indicating which predicates are satisfied by the object in grounded action ($|P|$ bits)
Global Features: The global node $g$ is initialized with a zero vector in $\mathbb{R}^{h}$ where $h$ is the chosen hidden representation dimension. This node can be used to aggregate and propagate graph-level information during message passing. This global node is essential in passing information at each round from nodes that are far apart. Hence, while nodes get their location information from neighboring edges, they also receive some global context at each round of the GNN ensuring that the number of GNN rounds does not become a bottleneck in passing information in large graphs. This is important since, as problems get larger the graph size grows, but the number of GNN rounds stays constant. The global node ensures all nodes and edges in the graph have some global context at each step.
@@ 3.3 Neural Architecture
Our neural architecture processes this graph representation through multiple components designed to handle the challenges of processing variable-sized inputs, capturing long range dependencies between objects and actions, and making sequential decisions to select actions and their arguments as shown in algorithm 1. We detail each component below:
3.3.1 GNN Encoder
@@
Graph Neural Networks (GNNs) (gnn_base, ) are particularly well-suited for encoding states in planning problems as they can naturally process relational structures while being invariant to permutations and handling varying input sizes (gnn_powerful, ). This allows them to learn patterns that generalize across different problem instances within the same domain, regardless of the number of objects involved. The key insight is that planning states are inherently relational - objects interact through predicates and actions - and GNNs can capture these relationships through message passing between nodes and edges. The GNN encoder (Algorithm 1, left side) transforms the input graph into learned embeddings.
Initialization (line 1): We initialize node features, edge features, and a global feature vector from the input graph. Message Passing (lines 3-9): For $L$ iterations, the encoder performs the following three update steps. 1) Edge Update (line 5): Each edge updates its representation by combining features from its connected elements. 2) Node Update (lines 7-8): Each node aggregates information from its neighboring edges along with global context. 3) Global Update (lines 9-10): The global vector is updated by aggregating information from all nodes and edges.
@@
The aggregation function (AGG) employs attention mechanisms to weight different elements according to their relevance. This architecture makes several key design choices motivated by the planning domain. First, we explicitly model and update edge representations because edges in our graph capture crucial information about action applicability. This edge information helps guide the model toward selecting valid and effective actions. Second, we include a global node that can rapidly propagate information across the graph. This is important because without the global node, information would need to flow through many message-passing steps to reach distant parts of the graph. The global node acts as a shortcut, allowing the model to maintain a comprehensive view of the planning state even as the number of objects and relations grows. After $L$ rounds of message passing have been performed, the final global node embedding $g^{l+1}$ captures the relevant planning context needed for action selection.
@@
1: Initialize: $\{\mathbf{v}^{0}_{i}\}_{i\in\mathcal{V}}$, $\{\mathbf{e}^{0}_{ij}\}_{(i,j)\in\mathcal{E}}$, $\mathbf{g}^{0}$
[…]
1: $h_{1}=\text{GRU}(\mathbf{g}^{L},\mathbf{0})$
2: // Select action schema 3: $\text{score}(a)=\text{MLP}(h_{1}\odot\mathbf{v}^{L}_{a})$ for each $a\in A$ 4: $a^{*}=\arg\max_{a\in A}\text{softmax}(\text{score}(a))$ 5: $h_{2}=\text{GRU}(h_{1},\mathbf{v}^{L}_{a^{*}})$ 6: // Select parameters 7: for $i=1$ to required parameter count do 8:   $\text{score}(o)=\text{MLP}(h_{i+1}\odot\mathbf{v}^{L}_{o})$ for each $o\in O$ 9:   $o^{*}_{i}=\arg\max_{o\in O}\text{softmax}(\text{score}(o))$ 10:   $h_{i+2}=\text{GRU}(h_{i+1},\mathbf{v}^{L}_{o^{*}_{i}})$ 11: return $(a^{*},o^{*}_{1},o^{*}_{2},\ldots,o^{*}_{n})$
Algorithm 1 Graph Attention-Based Action Ranking (GABAR)
3.3.2 GRU Decoder
The GRU-based decoder (Algorithm 1, right side) uses the encoded graph to select an action and its parameters. Initialization (line 1): The decoder initializes its hidden state using the global graph embedding and a zero vector. Action Schema Selection (lines 3-5): The decoder computes a score for each potential action schema (line 3). It then Selects the highest-scoring action (line 4) and updates the hidden state with the selected action’s features (line 5).
Parameter Selection (lines 7-10): For each required parameter position, scores are computed for all objects using the current hidden state (line 8). The highest-scoring object is selected as the parameter (line 9). The hidden state is updated with the selected object’s features (line 10). Output (line 11): The decoder returns the complete action grounding.
This autoregressive approach ensures that each parameter selection is conditioned on the graph structure, the selected action schema, and all previously selected parameters. By updating the GRU hidden state with each selection, the model captures dependencies between different components of the action. We repeat this greedily until all action parameters have been extracted. Since greedy parameter selection is often too myopic, we employ beam search to explore multiple choices in parallel. For a beam width $k$, at each step, we maintain the $k$ highest-scoring partial sequences. The final output is a ranked list of $k$ action groundings $(a,o_{1},\ldots,o_{n})$ along with their accumulated scores. This process of GRU-based decoding helps generalize the framework to handle actions with any arity, as we can extract as many objects as necessary for the action selected in the first step of the decoding. Hence, at any given state, we can use the decoder to extract the ranking over the set of actions with different arity all at once. This aligns with our goal of learning to rank fully grounded actions rather than independently optimizing each argument (extended beam search version of the decoder that maintains multiple candidate action groundings in parallel is provided as algorithm 2 in the Appendix). While training optimizes for selecting the optimal action, the learned model provides a ranking of the top $k$ grounded actions during execution. The planner then selects the first legal action in the current state according to the ranking. It also uses the ranking to avoid cycles - if an action leads to a state previously visited, then the next best action is considered.
3.4 Data Generation and Training